Fermi National Laboratory
 

Operation Summary for December 18, 2009 - December 21, 2009

 

The Integrated Luminosity for the period from 12/14/09 to 12/21/09 was 51.27 pb-1.
NuMI reported receiving 6.38E18 protons on target during this same period.

Friday Dec. 18
The day shift began with Operators preparing the Tevatron (TeV) for accesses, with Operations monitoring the stash (394.6E30) and the Antiproton Source stack (13.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments off due to the TeV and F-sector accesses, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments in access.

During the day shift, the TeV remained in access for maintenance work while the collision halls were in access for detector work. Operations reestablished beam to MiniBooNE by 1:30 PM and to Pbar and NuMI by 2:30 PM. Operations resumed sending beam to the MTest users at 3:30 PM.

During the evening shift, the TeV came out of access at 5 PM. Operations began the TeV turn-on. EE Support personnel changed out a corrector magnet power supply and then Operations began and completed an orbit check by 9:30 PM. Operations put the TeV into HEP shot setup at 10 PM and established store 7444 at 11:30 PM with an average initial luminosity (AIL) of 306E30. Operations disabled the MTest beam at 6 PM.

Saturday Dec. 19
The midnight shift began with Operations monitoring store 7444 (231.14E30) and the stash (185.7E10), with Pbar stacking (25.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations began sending beam to MTest at 4 AM.

During the day shift, Operations terminated store 7444 at 10:15 AM. Operations established store 7445 at 11:45 AM with an AIL of 274E30. At 12:30 PM, Operations reported that a vacuum leak had developed on a Booster RF cavity. RF personnel investigated and then with the help of Mechanical Support personnel began removing the RF cavity; this lasted the rest of the shift.

During the evening shift, RF and Mechanical Support personnel continued work on BRF19. Operations resumed sending beam to Pbar and NuMI at 11 PM. Operations reported at 11:30 PM that RF and Mechanical Support personnel had removed the Booster cavity and repaired the vacuum leak.

Sunday Dec. 20
The midnight shift began with Operations monitoring store 7445 and the stack, and with NuMI taking beam.

During the midnight shift, Operations resumed sending beam to MiniBooNE at midnight. Operations resumed sending beam to MTest at 4:15 AM.

During the day shifts, Operations terminated store 7445 at 9:30 AM and established store 7446 at 11 AM with an AIL of 258.3E30.

During the evening shift, at 9:35 PM, the Booster West Anode Power Supply (WAPS) began tripping off. Operations investigated and reported that this was due to a Booster RF station (BRF15).

Monday Dec. 21
The midnight shift began with Operations monitoring store 7446 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that the Booster WAPS continued to trip off throughout the shift. Operations terminated store 7446 at 3 AM and established store 7447 at 4:45 AM with an AIL of 264.6E30. Operations resumed sending beam to MTest at 5 AM. At 5:30 AM, a NuMI bulk power supply for the extraction line trim magnets failed. Investigation and repair of the power supply lasted until 7:45 AM.

The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator will allow general tuning today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/22/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory